# Multi-view Structural Convolution Network for Domain-Invariant Point Cloud Recognition of Autonomous Vehicles

### v3

@@ 3.2 Structural Aggregation Layer (SAL)
where $P_{in}\in\mathbb{R}^{N_{i-1}\times 3}$,$P_{out}\in\mathbb{R}^{N_{i}\times 3}$, $F_{in}\in\mathbb{R}^{N_{i-1}\times D_{i-1}}$, $F_{out}\in\mathbb{R}^{N_{i}\times D_{i-1}}$, and $N_{i}=\frac{N_{i-1}}{r}$. This local feature aggregation operation allows us to learn multi-scale local context features, which are crucial factors for extracting local geometric information.
@@
where $P_{in}\in\mathbb{R}^{N_{i}\times 3}$, $F_{in}\in\mathbb{R}^{N_{i}\times D_{i}}$, and $F_{out}\in\mathbb{R}^{D_{i}}$. By concatenating the global view feature with each row of output features from SCL$(D_{i-1},D_{i})$, we can obtain multi-view features $F\in\mathbb{R}^{N_{i}\times 2D_{i}}$. This concatenation results in SAL($D_{i-1},2D_{i}$), the output of this layer, and serves as the input to the next structural convolution layer.
3.3 Classification Process
@@
4 PROGRESSIVE DOMAIN EXPANSION PROCESS
In section 3, we explained how to obtain domain-invariant feature representations of point clouds by capturing geometric relationships between points. In this section, we focus on effectively generating unseen domains of 3d point clouds from specific domain point clouds, enabling MSCN to learn more robust feature representations.
Figure 3: The architecture for generating unseen domains in MSCN to learn more robust feature representations includes the following components. In (a), the modified MSCN is shown, consisting of a feature extractor (F), a classification head (C), and a projection head (P) with the projection head added for contrastive learning. In (b), the process of generating arbitrary unseen domains and training the MSCN with both source and generated data is illustrated. These generation and training processes are performed alternately.
4.1 Generator composition
@@
4.2 MSCN training with $\{S^{{}^{\prime}}\}^{k-1}_{i=1}$
@@
$L_{nce}(z_{i},z^{+}_{i})$ $=\min\limits_{F,C}-\log\left(\frac{\exp(z_{i}\cdot z^{+}_{i})}{\sum^{2N}_{j=1,j\neq i}\exp(z_{i}\cdot z_{j})}\right)$
@@
$L_{src}=L_{ce}(x_{i},y_{i})+L_{nce}(z_{i},z^{+}_{i})$ (13)
4.3 Unseen Domain $S^{{}^{\prime}}_{k}$ Generation
@@
$L_{gen,ce}=\min\limits_{G_{k},F,C}L_{ce}(C(F(G_{k}(x,n))),y),n\sim N(0,1)$ (15)
@@
5 EVALUATION
To evaluate the domain-invariant capabilities of the proposed MSCN, we conducted comprehensive experiments using both synthetic datasets and real-world datasets collected from self-driving vehicles. The synthetic dataset experiments aim to verify MSCN’s invariance to geometric transformations, including translation and scaling. These tests provide insights into the model’s ability to maintain performance under controlled yet challenging conditions.
In the real-world dataset experiments, we evaluated MSCN’s robustness in scenarios where the LiDAR resolution differs between training and testing environments. This setup simulates practical conditions in autonomous driving, where sensor configurations may vary. Furthermore, we assessed MSCN’s performance under severe domain shifts, combining both dataset-specific variations and additional transformations such as rotation and translation. Even in these harsh conditions, our MSCN consistently demonstrates robust performance, highlighting its effectiveness in domain-adaptive tasks.
5.1 Datasets & Implementation Detail
5.1.1 Synthetic Datasets
We conducted experiments using ShapeNetPart [37] and ModelNet40 [27] datasets. For the ShapeNetPart dataset, we evaluated the proposed MSCN on the part segmentation task, while for the ModelNet40 dataset, we performed object classification. Both experiments followed the standard training settings commonly used in previous work. In addition, specific experiments are conducted comparing MSCN with PointMLP [4] and PointTransformer [36] to demonstrate MSCN’s invariance to geometric transformations such as translation and scaling. For these experiments, we fixed the number of training epochs for all models to 30 and used a consistent point cloud size of 1024 points. These settings ensured a fair comparison across models and highlighted MSCN’s robustness under varying geometric transformations.
5.1.2 Real-world Indoor Dataset
5.1.3 Real-World Self-Driving Vehicle Datasets
We conducted experiments using three real-world point cloud datasets, namely KITTI[25], PanKyo[30], and nuScenes[26], as well as a synthetic dataset from the MORAI simulator[31]. These datasets, accessible through the following link  [^2] https://sites.google.com/site/cvsmlee/dataset, originate from various countries and manufacturers, making them crucial for evaluating the inference capacity of the model and the robustness to domain shifts.
Comparisons between our methods are conducted, including MSCN and MSCN, and precious methods including PointNet[1], PointNeXt[29], PointMLP[4], DGCNN[3], GCN[5], SCN[30], PointTransformer[36], and PointMamba[35]. PointMamba was first pre-trained on ShapeNet[37] and then trained on KITTI, PanKyo, and nuScenes for the classification task. The newtworks were evalauted through the a standard accuracy metric.
@@ 5.1.3 Real-World Self-Driving Vehicle Datasets
Table 1: Object part segmentation results on ShapeNetPart[21], grouped by dataset. mIOU values are reported for each method and dataset.
Figure 4: Examples of geometric transformation on ShapeNetPart[21], where scale 10 represents a transformation where the original point cloud is scaled by a factor of 10, while shift 90 indicates a random translation of the point cloud by a distance of 90 units.
5.2 Evaluation on Synthetic Datasets: Robustness to geometric deformation
We first conducted experiments on the ShapeNetPart [21] dataset for part segmentation. Table 1 summarizes the part segmentation results, where mIoU values are reported for each method. Our MSCN achieves performance comparable to state-of-the-art (SOTA) models, despite using only 1024 points, unlike other SOTA models that rely on 2048 points. While SOTA models exhibit high performance on the given dataset, MSCN’s performance is slightly inferior compared to SOTA models. This performance gap can be explained by the inherent trade-off between robustness and accuracy, as discussed in [57]. Our MSCN were designed to generalize well to harsh or out-of-distribution scenarios while preserving accuracy on natural data. However, Zhang et.al[57] showed this approach often leads to a shift in focus from minimizing standard classification errors (natural error) to improving robustness, which can degrade accuracy in ideal.
To further evaluate robustness of MSCN, we conducted experiments to visualize the robustness of MSCN compared to PointMLP and PointTransformer under geometric transformations. Figure 4 illustrates the results of scaling and translating point clouds. MSCN demonstrates robust part segmentation even under significant geometrical transformations, whereas PointMLP and PointTransformer show sensitivity to these changes, resulting in suboptimal performance. These results emphasize MSCN’s ability to generalize effectively across various geometric variations.
@@ 5.2 Evaluation on Synthetic Datasets: Robustness to geometric deformation
Table 2: Classification results on ModelNet40 [27].
Figure 5: Evaluation of Geometrical Invariance Properties on ModelNet40[27]. Each graph illustrates the accuracy of various models under specific geometric transformations applied to point clouds. (a) Translation: Accuracy as the translation magnitude increases sequentially from 0 to 200 units. (b) Scaling: Accuracy as the scale factor increases sequentially from 0.01 to 100.
Furthermore, we conducted experiments on the ModelNet40 [27] dataset for object classification. Table 2 presents the classification results, where MSCN is slightly lower than some state-of-the-art (SOTA) models. Although SOTA models demonstrate strong performance on this dataset, their sensitivity to geometric transformations can result in significant performance degradation in more challenging scenarios. To further investigate this, Figure 5 illustrates the performance of models under geometric transformations such as translation and scaling. MSCN consistently exhibits robust performance across all transformations, maintaining stable accuracy despite severe geometric changes. In contrast, other methods experience substantial drops in accuracy, highlighting their vulnerability to such transformations. These results underscore the strength of MSCN in terms of geometrical invariance, demonstrating its potential for real-world applications where robustness to geometric variations is critical.
However, it is important to note that ModelNet40 and ShapeNetPart are derived from CAD models, which are generated in highly idealized environments. In contrast, real-world point clouds are often sparse and only capture surface-level details, which makes them fundamentally different from synthetic datasets. To address this disparity, we extended our experiments to real-world point clouds, enabling a more comprehensive evaluation of our MSCN’s performance in practical scenarios.
@@
Table 3: Classification results on the ScanObjectNN[56] dataset. This table summarizes the performance of various methods under different experimental settings, including self-supervised pretraining, supervised learning, and the evaluation of MSCN’s translation and scale invariance. PT indicates Pretrained models, RT represents Random Translation, and RS denotes Random Scaling.
5.3 Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
We conducted experiments on the ScanObjectNN [56] dataset to evaluate the performance of MSCN on real-world indoor point clouds. Table 3 summarizes the classification results, demonstrating that MSCN achieves performance comparable to the state-of-the-art (SOTA) models in standard evaluation settings. Furthermore, MSCN exhibits remarkable robustness under geometric transformations such as scaling and translation, maintaining stable performance and demonstrating its strength in handling these variations.
However, it is important to note that ScanObjectNN consists of point clouds captured in indoor environments, which are relatively dense and structured. In contrast, point clouds captured in outdoor environments, such as those used in autonomous driving, are significantly sparser and often acquired under harsh conditions, including occlusions and uneven sampling. To address these challenges, we extended our experiments to self-driving vehicle datasets, enabling a more comprehensive evaluation of MSCN’s robustness in real-world scenarios.
@@ 5.3 Evaluation on Indoor Real-World Datasets: Comparable Performance and Robustness
Table 4: Intra-domain classification results on KITTI [25], nuScenes [26], and PanKyo [30] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
5.4 Evaluation on Self-Driving Datasets: Cross-Domain Robustness
@@
Our methods focus on maintaining consistent object recognition performance regardless of the type of point cloud data encountered. This means they are designed to work well not only with the data they were originally trained on but also with data from different sources. To thoroughly assess their effectiveness, we conducted two types of evaluations. Both intra-domain and cross-domain performances were evaluated to provide a comprehensive understanding of the capabilities of the methods and to ensure their reliability and effectiveness across various datasets.
@@
First, the performances are demonstrated within the KITTI[25], nuScenes[26], and PanKyo[30] datasets to capture classification performance seamlessly using the training dataset on the test dataset. Table 4 shows the classification performance within a single domain. These results measure the model’s ability to classify point clouds within the same domain used for training. While MSCN achieves competitive results, it lags slightly behind the SOTA methods in terms of accuracy. For example, the average intra-domain accuracy of MSCN across all datasets is 94.7%, compared to PointTransformer’s 98%. This performance gap can be explained from the bias-variance trade-off [38]. MSCN, designed for robust domain-invariant feature extraction, prioritizes reducing model variance to generalize effectively across diverse domains. As a result, its intra-domain performance exhibits slightly higher bias, which is a common trade-off observed in models designed with high generalization capacity. Specifically, while other SOTA models may minimize bias within the training domain, they often exhibit poor performance in domain-shift scenarios due to higher variance.
KITTI (64CH) PanKyo (128CH) nuScenes (32CH) Simulation (32CH) Average
@@
Table 5: Cross-domain classification results on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. The bold black, blue indicate the first and the second best performance on each scenario, respectively.
KITTI PanKyo nuScenes Simulation Average
Methods Param.(M) KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes Time(ms)
@@
Table 6: Inference time per point cloud on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. All values are measured in milliseconds (ms).
To demonstrate MSCN’s robustness, we further evaluated its performance under cross-domain scenarios as illustrated in Table 5, where substantial domain shifts between training and testing datasets. PointNet, PointNeXt, PointMLP, DGCNN, PointTransformer, and PointMamba demonstrate good performance when the domain shifts from simulation to the real world (Sim2Real) but show significantly lower performance when shifting from one real-world dataset to another (Real2Real). On the other hand, GCN and SCN show robust performance in Real2Real domain change scenarios but are sensitive in Sim2Real scenarios.
However, our MSCN demonstrates robust performance in both Real2Real and Sim2Real scenarios compared to the aforementioned methods. In particular, MSCN surpasses all SOTA methods with an average cross-domain accuracy of 82%, significantly higher than PointTransformer’s 66.2%. This 15.8% improvement highlights the effectiveness of MSCN in handling domain shifts, which is critical for real-world applications such as autonomous driving and robotics.
Furthermore, Table 6 presents the inference time required to predict the class of a single point cloud in both intra-domain and cross-domain settings. Models such as PointMLP, PointNeXt, and PointTransformer exhibit higher inference times due to their significantly large number of parameters. In contrast, although PointMamba is also Transformer-based and has a large parameter count, its inference time is relatively low. This efficiency can be attributed to its design, which likely focuses on reducing FLOPs (Floating Point Operations Per Second) to enhance computational efficiency.
In particular, the average inference time of MSCN is only 3.8 ms, which is well suited for real-world applications such as autonomous driving. Modern rotary LiDAR sensors typically operate at speeds of 10 Hz to 20 Hz, meaning one full rotation takes 0.05 to 0.1 seconds. Given that MSCN can process a single point cloud within 3.8 ms, it is capable of handling multiple scans within a single rotation, ensuring real-time performance without bottlenecking the data pipeline. This efficient processing time, combined with robust cross-domain performance, demonstrates the viability of MSCN for deployment in real-world autonomous systems.
Traning Dataset Test Dataset Layer Configuration Acc.(%)
KITTI PanKyo only SCL 78.9
@@
PanKyo KITTI only SCL 84.9
@@
nuScenes KITTI only SCL 92.1
@@
Table 7: Layer configuration evaluation results on KITTI [25], nuScenes [26], PanKyo [30], and simulation [31] datasets. The bold black indicates the best performance on each scenario.
5.5 Ablation study
In this section, we evaluate the compatibility of layer configurations of MSCN and the performance enhancement when MSCN was trained with synthetic point cloud datasets from our unseen domain generation method. In addition, the invariance properties of MSCN and MSCN are evaluated by conducting experiments on point cloud rotation and shift, further analyzing its robustness across varying domains.
Firstly, to assess the compatibility of layer configurations of MSCN, we compare the cases where there are configurations of both SCL and SAL, and only SCL. The former consists of four SCL and two SAL, which is illustrated in Figure 2 (a). The latter is composed of five SCL that are SCL(3, 32), SCL(32, 64), SCL(64, 128), SCL(128, 256), and SCL(256, 1024). Table 7 shows the performance of the methods in layer configurations where domain shift scenarios. MSCN outperforms the case that has only SCL without SAL during all domain shift scenarios, with an average performance that is 7.2% higher than the case that only consist of SCL.
Traning Dataset Test Dataset Method Acc.(%) Time (ms)
@@ 5.5 Ablation study
KITTI PanKyo MSCN 81.7 3.6
@@
PanKyo PanKyo MSCN 94.5 3.6
@@
nuScenes PanKyo MSCN 84.8 3.6
@@
Simulation PanKyo MSCN 83.9 3.6
@@
[…]

### v4

@@ III-B Structural Aggregation Layer (SAL)
where $P_{in}\in\mathbb{R}^{N_{i-1}\times 3}$$P_{out}\in\mathbb{R}^{N_{i}\times 3}$, $F_{in}\in\mathbb{R}^{N_{i-1}\times D_{i-1}}$, $F_{out}\in\mathbb{R}^{N_{i}\times D_{i-1}}$, and $N_{i}=\frac{N_{i-1}}{r}$. This local feature aggregation operation allows us to learn multi-scale local context features, which are crucial factors for extracting local geometric information.
@@
[…]
III-C Classification Process
@@
IV PROGRESSIVE DOMAIN EXPANSION PROCESS
In section III, we explained how to obtain domain-invariant feature representations of point clouds by capturing geometric relationships between points. In this section, we focus on effectively generating unseen domains of 3d point clouds from specific domain point clouds, enabling MSCN to learn more robust feature representations.
IV-A Generator Composition
@@
Figure 3: Progressive domain expansion process, including (a) the modified MSCN for contrastive learning and (b) the training architecture with the progressive domain expansion process.
IV-B MSCN Training with $\{S^{{}^{\prime}}\}^{k-1}_{i=1}$
@@
[…]
@@
[…]
IV-C Unseen Domain $S^{{}^{\prime}}_{k}$ Generation
@@
$\hskip-8.00003ptL_{\text{gen, ce}}=\min\limits_{G_{k},F,C}L_{ce}\!\left(C(F(G_{k}(x,n))),y\right),n\sim N(0,1)$ (15)
@@
V EVALUATION
To evaluate the domain-invariant capabilities of the proposed MSCN, we conducted comprehensive experiments using both synthetic datasets and real-world datasets collected from autonomous vehicles. The synthetic dataset experiments aim to verify MSCN’s invariance to geometric transformations, including translation and scaling. These tests provide insights into the model’s ability to maintain performance under controlled yet challenging conditions.
In the real-world dataset experiments, we evaluated MSCN’s robustness in scenarios where the LiDAR resolution differs between training and testing environments. This setup simulates practical conditions in autonomous driving, where sensor configuration may vary. Furthermore, we assessed MSCN’s performance under severe domain shifts, combining both dataset-specific variations and additional transformations such as rotation and translation. Even in these harsh conditions, our MSCN consistently demonstrates robust performance, highlighting its effectiveness in domain-adaptive tasks.
V-A Dataset and Implementation Details
V-A1 Synthetic Datasets
We conduct experiments using ShapeNetPart [48] and ModelNet40 [38] datasets. For the ShapeNetPart dataset, we evaluate the proposed MSCN on the part segmentation task, while for the ModelNet40 dataset, we perform object classification. Both experiments followed the standard training settings commonly used in previous work. In addition, specific experiments are conducted comparing MSCN with PointMLP [15] and PointTransformer [47] to demonstrate MSCN’s invariance to geometric transformations such as translation and scaling. For these experiments, we fixed the number of training epochs for all models to 30 and used a consistent point cloud size of 1024 points. These settings ensure a fair comparison across models and highlight MSCN’s robustness under varying geometric transformations.
Figure 4: Qualitative examples of geometric transformation on ShapeNetPart[32], where scale 10 represents a transformation where the original point cloud is scaled by a factor of 10, while shift 90 indicates a random translation of the point cloud by a distance of 90 units.
TABLE I: Object part segmentation results on ShapeNetPart[32]. mIOU values are reported for each method.
@@ V-A1 Synthetic Datasets
V-A2 Real-world Dataset
V-A3 Real-World Autonomous Vehicle Datasets
We conduct experiments using three real-world point cloud datasets, namely KITTI[36], PanKyo[41], and nuScenes[37], as well as a synthetic dataset from the MORAI simulator[42]. These datasets, accessible through the following link  [^1] https://sites.google.com/site/cvsmlee/dataset, originate from various countries and manufacturers, making them crucial for evaluating the inference capacity of the model and the robustness to domain shifts.
Comparisons between our methods are conducted, including MSCN and MSCN, and precious methods including PointNet[12], PointNeXt[40], PointMLP[15], DGCNN[14], GCN[16], SCN[41], PointTransformer[47], and PointMamba[46]. PointMamba was first pre-trained on ShapeNet[48] and then trained on KITTI, PanKyo, and nuScenes for the classification task. The networks are evaluated through the a standard accuracy metric.
V-B Evaluation on Synthetic Datasets: Robustness to Geometric Deformation
We first conduct experiments on the ShapeNetPart [32] dataset for part segmentation. Table I summarizes the part segmentation results, where mIoU values are reported for each method. Our MSCN achieves performance comparable to state-of-the-art (SOTA) models, despite using only 1024 points, unlike other SOTA models that rely on 2048 points. While SOTA models exhibit high performance on the given dataset, MSCN’s performance is slightly inferior compared to SOTA models. This performance gap can be explained by the inherent trade-off between robustness and accuracy, as discussed in [68]. They showed that models performing robustly often lead to a shift in focus from minimizing standard classification errors (natural error) to improving robustness, which can degrade accuracy in the ideal condition (e.g. same properties between training and test dataset). Note that our MSCN were designed to generalize well to harsh or out-of-distribution scenarios while preserving accuracy on natural data. To evaluate the robustness of MSCN, we conduct experiments to visualize the robustness of MSCN compared to PointMLP and PointTransformer under geometric transformations. Figure 4 illustrates the results of scaling and translating point clouds. MSCN demonstrates robust part segmentation even under significant geometrical transformations, whereas PointMLP and PointTransformer show sensitivity to these changes, resulting in suboptimal performance. These results emphasize MSCN’s ability to generalize effectively across various geometric variations.
Figure 5: Evaluation of geometrical invariance properties on ModelNet40 [38], showing model accuracies under translation (0–200 units) and scaling (0.01–100).
TABLE II: Classification results on ModelNet40 [38].
@@ V-B Evaluation on Synthetic Datasets: Robustness to Geometric Deformation
Furthermore, we conduct experiments on the ModelNet40 [38] dataset for object classification. Table II presents the classification results, where MSCN is slightly lower than some SOTA models. Although SOTA models demonstrate strong performance under ideal conditions, their sensitivity to geometric transformations can result in significant performance degradation in more challenging scenarios. To further investigate this, Figure 5 illustrates the performance of models under geometric transformations such as translation and scaling. MSCN consistently exhibits robust performance across all transformations, maintaining stable accuracy despite severe geometric changes. In contrast, other methods experience substantial drops in accuracy, highlighting their vulnerability to such transformations. These results underscore the strength of MSCN in terms of geometrical invariance, demonstrating its potential for real-world applications where robustness to geometric variations is critical.
However, it is important to note that ModelNet40 and ShapeNetPart are derived from Computer Aided Design (CAD), which are generated in highly idealized environments. In contrast, real-world point clouds are often sparse and only capture surface-level details, which makes them fundamentally different from synthetic datasets. To address this disparity, we extend our experiments to real-world point clouds, enabling a more comprehensive evaluation of our MSCN’s performance in practical scenarios.
TABLE III: Classification results on the ScanObjectNN[67] dataset. PT indicates Pretrained models, RT represents Random Translation, and RS denotes Random Scaling.
@@
TABLE IV: Intra-domain classification results on AV datasets.
@@
V-C Evaluation on Real-World Datasets: Comparable Performance and Robustness
We conduct experiments on the ScanObjectNN [67] dataset to evaluate the performance of MSCN on real-world indoor point clouds. Table III summarizes the classification results, demonstrating that MSCN achieves performance comparable to the SOTA models in standard evaluation settings. Furthermore, MSCN exhibits remarkable robustness under geometric transformations such as scaling and translation, maintaining stable performance and demonstrating its strength in handling these variations.
However, it is important to note that ScanObjectNN consists of point clouds captured in indoor environments, which are relatively dense and structured. In contrast, point clouds captured in outdoor environments, particularly used in autonomous driving, are significantly sparser and often acquired under harsh conditions, including occlusions and uneven sampling. To show MSCN’s effectiveness in these challenges, we extend our experiments to AV datasets.
V-D Evaluation on AV Datasets: Robustness to Different Sensor Configurations and Geometric Deformations
@@
Our methods focus on maintaining consistent object recognition performance regardless of the type of point cloud data encountered. This means MSCN was designed to work well not only with the data they were originally trained on but also with data from different sources. To thoroughly assess their effectiveness in the AV dataset, we conduct two types of evaluations. Both intra-domain and cross-domain performances are evaluated to provide a comprehensive understanding of the capabilities of the MSCN and to ensure its reliability and effectiveness across various AV datasets.
@@
Table IV shows the classification performance within a single domain. These results measure the model’s ability to classify point clouds within the same domain used for training. While MSCN achieves competitive results, it lags slightly behind the SOTA methods in terms of accuracy. This performance gap can be explained from the bias-variance trade-off [49]. MSCN, designed for robust domain-invariant feature extraction, prioritizes reducing model variance to generalize effectively across diverse domains. As a result, its intra-domain performance exhibits slightly higher bias, which is a common trade-off observed in models designed with high generalization capacity. Specifically, while other SOTA models may minimize bias within the training domain, they often exhibit poor performance in domain-shift scenarios due to higher variance.
To investigate SOTA models’ sensitivity and MSCN’s robustness under the different LiDAR configurations, we further evaluate its performance under cross-domain scenarios as illustrated in Table V, where substantial domain shifts between training and testing datasets. PointNet, PointNeXt, PointMLP, DGCNN, PointTransformer, and PointMamba demonstrate good performance when the domain shifts from simulation to the real world (Sim2Real) but show significant performance degradation when shifting from one real-world dataset to another (Real2Real). This is because real-world datasets not only differ in resolution but also exhibit harsh characteristics such as irregular sampling, occlusion, and sensor noise. These models tend to overfit to specific point cloud properties, leading to substantial recognition failures when deployed across diverse real-world LiDARs, which poses a critical risk for safety-critical applications in AVs. On the other hand, GCN and SCN show robust performance in Real2Real domain change scenarios but are sensitive in Sim2Real scenarios. This sensitivity arises because their graph-based structures adapt well to irregular real-world point clouds but struggle to generalize from the uniform and noise-free distribution of simulated point clouds. Consequently, their limited ability to bridge the Sim2Real gap hinders the use of large-scale simulation data for training, thereby restricting scalability and reducing the cost-effectiveness of model development.
However, our MSCN demonstrates robust performance in both Real2Real and Sim2Real scenarios compared to the aforementioned methods. In particular, MSCN surpasses all SOTA methods with an average cross-domain accuracy of 82%, significantly higher than PointTransformer’s 66.2%. This 15.8% improvement highlights the effectiveness of MSCN in handling domain shifts, which is critical for the deployment of AVs.
Furthermore, Table VI presents the inference time required to predict the class of a single point cloud in both intra-domain and cross-domain settings. Models such as PointMLP, PointNeXt, and PointTransformer exhibit higher inference times due to their significantly large number of parameters. In contrast, although PointMamba is also Transformer-based and has a large parameter count, its inference time is relatively low. This efficiency can be attributed to its design, which likely focuses on reducing FLOPs (Floating Point Operations Per Second) to enhance computational efficiency.
In particular, the average inference time of MSCN is only 3.8 ms, which is well suited for AVs. Modern rotary LiDAR sensors typically operate at speeds of 10 Hz to 20 Hz, meaning one full rotation takes 0.05 to 0.1 seconds. Given that MSCN can process a single point cloud within 3.8 ms, it is capable of handling multiple scans within a single rotation, ensuring real-time performance without bottlenecking the data pipeline. This efficient processing time, combined with robust cross-domain performance, demonstrates the viability of MSCN for deployment in AV systems.
Finally, to further reflect the operational conditions of real-world AVs, we evaluate the robustness of our model not only under sensor configuration changes but also under continuous geometric variations in point clouds that naturally occur during driving. As AVs move, the surrounding point clouds are constantly subjected to shifts and rotations due to changes in viewpoint and vehicle trajectory. To replicate this, we progressively rotate the point clouds from 0 to 100 in 10 increments and randomly shift them within a range of 0 to 100 units, while simultaneously considering cross-domain sensor configuration changes.
Figure 6 presents six different scenarios that combine rotation and shift transformations in both intra-domain and cross-domain settings. In the intra-domain experiments, most baseline models suffer a sharp decline in accuracy as the rotation or shift magnitude increases, whereas our MSCN and MSCN show only a gradual decrease, demonstrating strong resilience to geometric perturbations. More importantly, in cross-domain scenarios such as PanKyo $\rightarrow$ KITTI and PanKyo $\rightarrow$ nuScenes, our models consistently maintain superior accuracy under substantial rotations and shifts, outperforming all baselines. These results confirm that MSCN and MSCN are not only effective at handling domain shifts across heterogeneous LiDAR sensors, but are also robust to dynamic geometric transformations, making them well-suited for deployment in the challenging and continuously changing environments encountered by real-world AVs.
TABLE V: Cross-domain classification across both real-world AV and simulation datasets.
KITTI (64CH) [36] PanKyo (128CH) [41] nuScenes (32CH) [37] Simulation (32CH) [42] Average
@@
TABLE VI: Evaluation on Inference time across AV datasets. All values are measured in milliseconds (ms).
KITTI [36] PanKyo [41] nuScenes [37] Simulation [42] Avg.
Methods Param. KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes KITTI PanKyo nuScenes Time
@@
Figure 6: Robustness evaluation under rotations and shifts in both intra- and cross-domain settings.
TABLE VII: Layer configuration evaluation results on both real-world AV and simulation datasets.
Training Dataset Test Dataset Layer Configuration Acc.(%)
KITTI [36] PanKyo only SCL 78.9
@@
PanKyo [41] KITTI only SCL 84.9
@@
nuScenes [37] KITTI only SCL 92.1
@@
TABLE VIII: Unseen domain generation effects on both real-world AV and simulation datasets.
Training Dataset Test Dataset Method Acc.(%) Time (ms)
@@
KITTI [36] PanKyo MSCN 81.7 3.6
@@
Pankyo [41] PanKyo MSCN 94.5 3.6
@@
nuScenes [37] PanKyo MSCN 84.8 3.6
@@
Simulation [42] PanKyo MSCN 83.9 3.6
@@
V-E Ablation study
We evaluate the compatibility of layer configurations of MSCN and the performance enhancement when MSCN was trained with synthetic point cloud datasets from our unseen domain generation method.
First, to assess the compatibility of layer configurations of MSCN, we compare the cases where there are configurations of both SCL and SAL, and only SCL. The former consists of four SCL and two SAL, which is illustrated in Figure 2 (a). The latter is composed of five SCL that are SCL(3, 32), SCL(32, 64), SCL(64, 128), SCL(128, 256), and SCL(256, 1024). Table VII shows the performance of the methods in layer configurations where domain shift scenarios. MSCN outperforms the case that has only SCL without SAL during all domain shift scenarios, with an average performance that is 7.2% higher than the case that only consist of SCL.